### v1

# XPrompt: Explaining Large Language Model’s Generation via Joint Prompt Attribution

Abstract
Large Language Models (LLMs) have demonstrated impressive performances in complex text generation tasks. However, the contribution of the input prompt to the generated content still remains obscure to humans, underscoring the necessity of elucidating and explaining the causality between input and output pairs. Existing works for providing prompt-specific explanation often confine model output to be classification or next-word prediction. Few initial attempts aiming to explain the entire language generation often treat input prompt texts independently, ignoring their combinatorial effects on the follow-up generation. In this study, we introduce a counterfactual explanation framework based on joint prompt attribution, XPrompt, which aims to explain how a few prompt texts collaboratively influences the LLM’s complete generation. Particularly, we formulate the task of prompt attribution for generation interpretation as a combinatorial optimization problem, and introduce a probabilistic algorithm to search for the casual input combination in the discrete space. We define and utilize multiple metrics to evaluate the produced explanations, demonstrating both faithfulness and efficiency of our framework.
1 Introduction
Large Language Models (LLMs), such as GPT4 [1], LLaMA [2] and Claude [3], have shown excellent performance in various natural language generation tasks including question answering, document summarization, and many more. Despite the great success of LLMs, we still have very limited understanding on the LLM generation behavior – which parts in the input cause the model to generate a certain sequence. Unable to explain the causality between the input prompt and the output generation could cause failure in recognizing potential unintended consequences, such as harmful response [4, 5, 6, 7] and biased generation [8] attributed to a specific malicious description in the input. These issues undermine human trust in model usage, thus highlighting a pressing need for developing an interpretation tool that attributes how an input prompt leads to the generated content.
Explaining LLM generation through prompt attribution involves extracting the most influential prompt texts on the model’s entire generation procedure, a realm that remains relatively under-explored in current research. While extensive works on input attribution are proposed for text classification interpretation [9, 10] and next-word generation rationale [11, 12], they can not be directly applied to explain the full generation sequence due to its complicated joint probability landscape and the autoregressive generation procedure. The complexity for interpreting LLM generation compounds as the model size increases.
Another line of works involves prompting LLMs to self-explain their behaviors [13]. This method relies on the model’s innate reasoning capabilities, although current findings suggest that these capabilities may not always be faithful [14, 15].
There are limited existing attempts focusing on explaining the relationship between the input prompts and the complete generated sequence. The most relevant approach, Captum [16], sequentially determines the importance score for each token by calculating the variations in the joint probability of generating the targeted output sequence when the token is dropped from the model input. While being straightforard, this approach treats tokens as independent features, ignoring their joint semantic influence on the generated output. In fact, tokens may contain overlapping or complementary information. For example, given the input prompt: “Write a story about the doctor and his patient”, the most influential components are “doctor” and “patient”. Individually removing either of these words would not significantly alter the generated output, resulting in inaccurately low importance score for each of them. This is caused by the semantic interaction among these components, allowing the model to infer the meaning of the omitted one. Existing attribution methods that ablate each token individually fail to capture such combinatorial effect. A straightforward remedy might involve exhaustively assessing all possible combinations to observe the variations in the model generations, which however is impractical due to the vast search space with long-context input prompts.
To efficiently search the space for generating accurate prompt explanations, we develop our framework XPrompt, which provides the counterfactual explanation to highlight which components of input prompts have the fundamental effect on the generated context via solving a combinatorial optimization problem.
We aim to explain the generation behavior of model outputs for any given prompt while take the joint effects of the prompt components into account. Assuming that removing the essential parts of the prompt would result in a significant variation in the model’s output, we propose the novel objective function and formulate our task of providing faithful counterfactual explanations for the input prompt as an optimization problem. To quantify the influence of token combinations in the prompt on the generations, we incorporate a mask approach for joint prompt attribution. Thus, our goal of extracting the explanations has been converted to finding the optimal mask of the input prompt. We solve this problem by a probabilistic search algorithm, equipped with gradient guidance and probabilistic updaten for efficient exploration in the discrete solution space. Our main contributions could be summarized as follows:
  • We propose a general interpretation scheme for LLM generation task that attributes input prompts to the entire generation sequence. Notably, this recipe considers the joint influence of input token combinations on the generation. This motivation naturally formulates a combinatorial optimization problem for explaining generation with the most influential prompt texts.
  • We demonstrate XPrompt, an efficient probabilistic search algorithm to solve the optimization problem. XPrompt works by searching better token combinations that lead to larger generation changes. It takes the advantage of both the gradient information and the probabilistic search-space strategy, thereby achieving an efficient prompt interpretation tool.
  • Our framework demonstrates strong performance on language generation tasks including text summarization, question-answering, and general instruction datasets. The faithfulness of our explanations is evaluated based on a suit of comprehensive metrics considering generation probability, word frequency, and semantic similarity, verifying the transferability and effectiveness of our methods across a variety of tasks.
2 Related Work
While there are extensive works devoted to explaining language models in the context of text classification tasks [17, 18, 19], relatively few attempts [11, 12] are proposed to investigate the importance of prompt texts on the entire generation procedure especially for LLMs. This demonstrates a research gap that this work aims to fill in.
Explaining Language Generation
There are many work explaining the predictions generated by LLMs by measuring the importance of the input features to the model’s prediction on the classification tasks. One group of studies perturb the specific input by removing, masking, or altering the input features, and evaluate the model prediction changes [20, 21]. The other group of works, such as integrated gradients (IG) [22] and first-derivative saliency [23], leverage the gradients of the output with respect to the input to determine the input feature importance.
Moreover, a few studies utilize the surrogate model to explain the individual predictions of the black-box models, and the representative method is called LIME [24]. As for explaining the model’s generation behavior, Captum [16] calculates the token’s importance score by sequentially measuring the contribution of input token to the output, which lacks of the accounting for the semantic relationships between tokens. Another work, ReAGent [11], focuses on the next-word generation task, computing the importance distribution for the next token position. This method ignores the contextual dependencies in the the generated output and could not adequately account for dynamically changing generations. Our framework aims to interpret the joint effects of the input prompts on the entire output contexts with considering of the textual information covered the input prompt.
Self-explaining by Prompting
As language models increase in scale, prompting-based models demonstrate remarkable abilities in reasoning [25], creativity [26], and adaptability across a range of tasks [27]. However, the complex reasoning processes of these models remain elusive and require tailored paradigm to better understand the prompting mechanism. For instance, the chain-of-thought (CoT) paradigm could explain the LLM behaviors by prompting the model to generate the reasoning chain along with the answers [13], as pre-trained LLMs have demonstrated a certain ability to self-explain their behaviors. However, recent studies have also suggested that the reasoning chain does not guarantee faithful explanations of the model’s behavior [28] and the final answer might not always follow the generated reasoning chain [14]. xLLM [29] enhances the fidelity of explanation derived from LLMs via a fidelity evaluator, which however is designed for classification tasks. Our efforts are concentrated on explaining LLM generation by analyzing the attribution of the input prompts to the output content, without depending on the model’s innate reasoning ability that are not yet satisfactory.
3 Preliminaries
We first introduce necessary notions depicting the LLM generation process, and then discuss limitations of prior attempts explaining the entire generation via prompt attribution.
Notions on LLM Generation
Denote a specific input prompt as a sequence of tokens $\bm{x}=(x_{1},\dots,x_{T})$, $x_{i}\in\{1,2,\dots,|V|\}$, where $|V|$ represents the vocabulary size, $T$ is the length of the input sequence and the set of all token indices is =$\mathcal{I}=\{1,2,\dots,T\}$. The corresponding generated output $\bm{y}$ could be represented as a sequence of tokens $\bm{y}=(y_{1},\dots,y_{S})$ with $y_{j}\in\{1,2,\dots,|V|\}$. The output tokens $\{y_{i}\}_{i=1}^{S}$ are generated from the LLM $f_{\bm{\theta}}$ parameterized by $\bm{\theta}$ in an autoregressive manner with the probability $p_{\bm{\theta}}(\bm{y}|\bm{x})$ as:
$p_{\bm{\theta}}(\bm{y}|\bm{x})=p_{\bm{\theta}}(y_{1}|\bm{x})\prod_{i=1}^{S-1}p_{\bm{\theta}}(y_{i+1}|\bm{x},y_{i}).$ (1)
The probability of generating the output text $\bm{y}$ given the input prompt $\bm{x}$ illustrates that the coherence and generation of the output texts heavily rely on the input prompt $\bm{x}$, indicating an implicit causal relationships between the input prompt $\bm{x}$ and the output $\bm{y}$.
Limitation of Prior Interpretation Attempt
There are limited prior works focus on explaining the relationship between the individual input prompts and the entire generated sequence, and their faithfulness is limited by treating input tokens independently. Specifically, Captum [16], calculates the importance score for each token by slightly perturbing the input $\bm{x}$ at the $i$-th token:
$G_{i}(\bm{x};\bm{\theta})=p_{\bm{\theta}}(\bm{y}|\bm{x})-p_{\bm{\theta}}(\bm{y}|\bm{x}_{\mathcal{I}\setminus\{i\}}),$ (2)
where $\mathcal{I}$ denotes all token indices. Through Eq. (2), one can calculate an attribution score for each token $i$ indicating its importance towards the original generation result $\bm{y}$. Such method is direct and simple, however, it treats tokens as independent features, ignoring their semantic interaction and joint effect on the generated output. To make it more obvious, let’s recall the previous example of “Write a story about a doctor and his patient.”. Note that since “doctor” and “patient” are semantically correlated, and removing any of the two words would not have a significant influence on the generated response, their corresponding importance score will not be too high. On the contrary, the most important token detected by this approach would be “his”, which is clearly not ideal.
These observations inspire us to develop a new prompt attribution method that considers the semantic correlation among tokens. Specifically, we assume that the content generated by LLMs is primarily influenced by a subset of tokens jointly. The remaining tokens serve as auxiliary or potentially less relevant information. Those subset of tokens, which fundamentally shape the model’s output, are viewed as the explanatory tokens for the generated content. In other words, explanatory tokens do not affect the model output independently, but jointly contribute to the generated responses.
4 Proposed Method
In this section, we propose XPrompt, a simple yet effective framework designed for generative tasks to explain the attribution of the input prompt by solving a discrete optimization problem. We start with formulating the general objective for the discrete optimization problem, and then we introduce our proposed probabilistic search algorithm for solving the problem.
Figure 1: Overview of XPrompt. Left: Demonstrating the pipline of the algorithm. Right: Illustrating the process of mask $\bm{m}$ sampling.
Problem Formulation for Joint Prompt Attribution
In order to explain language generation via prompt attribution, instead of treating the prompt tokens independently as in previous works [16], we propose to evaluate the joint effect of $k$ prompt tokens on the generated sequence.
Specifically, consider a binary prompt mask $\bm{m}=(m_{1},\dots,m_{T})$ in which $m_{i}\in\{0,1\}$ indicates whether the $i$-th token is important or not. The joint probability of generating the original output sequence $\bm{y}$ given an input masked context $\bm{m}\odot\bm{x}$ is computed as $p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})$, where $\odot$ denotes the Hadamard product. In our paper, we aim to identify a binary mask $\bm{m}$ that maximizes the discrepancy in the probability of generating the same output $\bm{y}$ when comparing the masked input $\bm{m}\odot\bm{x}$ to the original input $\bm{x}$. A large discrepancy indicates that the masked token combination are the most influential components for generating $\bm{y}$, thus should jointly serve as the attributed explanation. Consequently, this involves training the binary mask $\bm{m}$ to optimize the following objective function:
$\max_{\bm{m}\in\{0,1\}^{T}}\mathcal{L}(\bm{m},\bm{x},\bm{y};\bm{\theta}):$ $=p_{\bm{\theta}}\left(\bm{y}|\bm{x}\right)-p_{\bm{\theta}}\left(\bm{y}|\bm{m}\odot\bm{x}\right)$ (3)
[…]
where $k$ represents the number of explanatory tokens. Intuitively, optimizing the objective outlined in Eq. (3) suggests that we want to find the top $k$ important tokens which, if they are masked, lead to a substantial variation in model’s output probabilities. Compared with prior methods for individual token attribution stated in Eq. (2), our formulation in Eq. (3) measures the joint attribution of a subset of tokens, capturing token interactions to enable more accurate prompt explanations.
Challenges in Solving Eq. (3)
Note that Eq. (3) is a constraint discrete optimization problem that is non-trivial to solve. One naive solution would be transforming the discrete optimization problem into a continuous one, i.e., let $\bm{m}\in[0,1]^{T}$, and formulate the constraint into a regularization term. Then one can adopt traditional gradient descent based optimization solutions to solve the problem. However, such a strategy would require extensive gradient calculations and backward steps on the original inputs, which can be inefficient in practice especially for LLMs. Moreover, the obtained continuous mask is not the final explanation we want. In fact, the approximation error when transforming the continuous mask back into the discrete space can also be quite significant, leading to worse performances. Another straightforward solution involves searching through all possible token combinations exhaustively. However, this could be impractical as well due to the enormous search space especially while facing long-context inputs. Therefore, we hope to develop a new method that adopts a search-based strategy to simplify the algorithm design and satisfy our constraints, while also leveraging gradient information for efficient optimization.
Algorithm 1 Explainable Prompt Generator: XPrompt
Input: Input tokens $\bm{x}$, output tokens $\bm{y}$, the integer $k$ denoting the number of explanatory tokens, $1\leq k\leq T$, input mask $\bm{m}^{(0)}=\textbf{1}$, and the sampling numbers $N$.
Output: Optimal mask $\bm{m}^{(N)}$
1:$\bm{g}=|\nabla_{\bm{m}^{(0)}}\mathcal{L}(\bm{m}^{(0)},\bm{x},\bm{y};\bm{\theta})|$
[…]
3:for $n=1$ to $N$ do
4:     Sample $l\sim\text{softmax}\big{(}\bm{m}^{(n)}\odot\bm{g}\big{)}$
5:     Sample $v\sim\text{softmax}\big{(}(\mathbf{1}-\bm{m}^{(n)})\odot\bm{g}\big{)}$
6:     $\bm{m}^{\text{tmp}}=\bm{m}^{(n)}.\text{copy}()$; switch the value of ${m}^{\text{tmp}}_{l}$ and ${m}^{\text{tmp}}_{v}$
7:     if $p_{\bm{\theta}}(\bm{y}|{\bm{m}^{\text{tmp}}}\odot\bm{x})<p_{\bm{\theta}}(\bm{y}|{\bm{m}^{(n)}}\odot\bm{x})$ then
8:         $\bm{m}^{(n+1)}=\bm{m}^{\text{tmp}}$
9:     else
10:         $\bm{m}^{(n+1)}=\bm{m}^{(n)}$
11:     end if
12:end for
Proposed Probabilistic Search Algorithm
To tackle this challenge, we propose XPrompt, a novel explainable prompt generator, for efficiently obtaining an optimal solution for Eq. (3). We summarize our XPrompt in Algorithm 1. The high-level pipeline is illustrated in Figure 1 left: we initialize and maintain exactly $k$ entries in the mask $\bm{m}$ to be zero to enforce the constraint and capture their joint influence of being masked; the mask is then iteratively updated by sampling indexes for value swapping, searching towards the direction with increased generation loss $\mathcal{L}(\bm{m},\bm{x},\bm{y};\bm{\theta})$. At the core of this pipeline is the sampling and update of the discrete mask, which demands an efficient exploration in the vast search space. Figure 1 right shows this step, which is featured by the following two essential components, the gradient-guided masking and the probabilistic search update.
Gradient-Guided Masking: Trivial solutions that set the mask $\bm{m}$ by uniformly random could cost massive sampling to hit the right optimization direction. Gradient is a common indicator of feature importance, as evidenced in existing practices [30, 6, 31]. Therefore, we propose to use gradient as a guidance to set and sample the mask for more efficient optimization. Specifically, we begin with the binary mask $\bm{m}^{(0)}=\bm{1}$ which indicates that all tokens in the input $\bm{x}$ are marked as non-explanatory ones. Then we compute the gradient $\nabla_{\bm{m}^{(0)}}\mathcal{L}(\bm{m}^{(0)},\bm{x},\bm{y};\bm{\theta})$ of the loss function in Eq. (3), and denote the magnitude of gradients as $\bm{g}=|\nabla_{\bm{m}^{(0)}}\mathcal{L}(\bm{m}^{(0)},\bm{x},\bm{y};\bm{\theta})|$. Note that the components with larger gradient magnitudes in $\bm{g}$ imply that altering the corresponding tokens could result in a sharper change to the generated outputs. In order to initialize the explanatory $k$ tokens guided by the gradient magnitude, we set the binary mask $\bm{m}^{(1)}$ where $m^{(1)}_{i}=0$ indicates $g_{i}$ is the top-$k$ value in $\bm{g}$. The gradient guidance and mask initialization are obtained following Line 1-2 in the algorithm.
Probabilistic Search Update: While gradient is informative, greedily determining explanatory tokens by top gradients lacks exploration, leading to suboptimal solutions. Therefore, we propose a probabilistic search mechanism for mask sampling and update. Specifically when updating the current binary mask $\bm{m}$, we iteratively sample a non-zero entry $l$ from the mask and swap its value with a sampled zero entry $v$ to explore a new (and potentially better) solution for the mask. Particularly, the sampling is also guided by the gradient calculated before: the non-zero entry in the mask (representing non-explanatory tokens) is sampled following probabilities calculated by the normalized gradient magnitudes, i.e., softmax($\bm{m}^{(n)}\odot\bm{g}$), and we employ a similar sampling strategy for zero entries (explanatory tokens). After swapping the mask indicators for the two sampled tokens $l$ and $v$, we generate a temporary mask $\bm{m}^{\text{tmp}}$. We then evaluate whether this temporary mask leads to a decrease in output probability. If it does, we update the binary mask to $\bm{m}^{\text{tmp}}$, otherwise we leave the binary mask as is. Consequently, without requiring intensive gradient computations, these sampling iterations keep discovering improved solutions for the discrete optimization problem shown in Eq. (3). We conclude this update by sampling process in Line 3-12.
While capturing the joint influence of being masked, the proposed XPrompt takes the advantage of both the gradient information and the search-space strategy, thereby achieving better efficiency than adopting either method alone. XPrompt requires only a single step of gradient calculation and avoids the need to convert between discrete and continuous masks. The obtained gradient information provides a favorable initial searching direction and reliable sampling probabilities that enhances the search efficiency.
5 Experiment
This section aims to verify the effectiveness and efficiency of our proposed framework for interpreting the LLMs on the generation task. We conduct the experiments to answer the following questions:
  • Q1: Do the generated prompt attributions play a predominant role in the model’s generation, thereby serving as faithful counterfactual explanations for the generated output?
  • Q2: Is our interpretation algorithm efficient for practical usage?
  • Q3: Could the proposed algorithm effectively identify a combination of tokens that impose joint influence on the model generation?
We provide quantitative studies to evaluate the faithfulness of the explanatory prompt fragments generated by XPrompt, comparing with existing interpretation baselines and ablation variants.
5.1 Experiment Settings
Models & Baselines
In the experiment, we employ two LLMs as our targeted $f_{\bm{\theta}}(\cdot)$: LLaMA2 (7B-Chat) [2] and Vicuna (7B) [32]. There are relatively few methods attributing prompts on the entire language generation, and we choose random removal (Random), Integrated-Gradient [22] and Captum [16] as our baseline for comparison. We implement these models using the PyTorch framework and pretrained weights from the transformers Python library [33], and conduct our experiments on an Nvidia RTX A6000-48GB platform with CUDA version 12.0.
Datasets
We employ three distinct text generation datasets: Alpaca [34], tldr_news [35], and MHC [36], to evaluate the effectiveness of our method across various generation tasks. As we aim to capture the joint influence of prompts on the model generations, we focus on relatively long-context prompts rather than simple one-sentence prompts. Longer prompts tend to contain more diverse vocabularies, convey more information, and thus more likely to exhibit high textual correlation.
For evaluation purposes, we randomly select approximately 110 data samples with at least 15 words from each dataset. All datasets are publicly available and more details about the dataset are illustrated in Appendix A.1.
5.2 Evaluation Metrics
Faithfulness scores are a key metric for assessing the quality of explanations, with faithful explanations accurately reflecting the model’s decision-making process [28]. Studies [37, 38] suggest that if a certain input tokens are truly important, their removal should lead to a more significant change in model output than the removal of random tokens [39]. Therefore, after removing explanatory tokens identified by a faithful method, the model’s new generation would significantly differ from using the original output. Moreover, the input prompt with masking these explanatory tokens are less likely to reproduce the original model response $\bm{y}$, indicating a lower value of $p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})$.
To quantitatively evaluate explanation faithfulness, we measure the change of model generation behavior from two dimensions. On one hand, we compare the model’s original and new generations: the originally generated sequence $\bm{y}$ is based on the complete input prompt $\bm{x}$ (e.g., $\bm{y}=f_{\bm{\theta}}(\bm{x})$), while the new generation is conditioned on the masked prompt (e.g., $\bm{y}^{\prime}=f_{\bm{\theta}}(\bm{m}\odot\bm{x})$).
We thus measure the similarity of the original generation $\bm{y}$ and the new generation $\bm{y}^{\prime}$ based on their word frequency and semantics. A smaller similarity reflects a larger change in model generation, suggesting a better explanation. On the other hand, we measure the likelihood of generating the original output $\bm{y}$ when the model uses the masked prompt, e.g., changes on $p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})$. A smaller likelihood indicates better explanations whose mask prevents the model from generating its original output. Detailed definitions of these metrics are explained below.
Word Frequency: BLEU [40] is widely used to measure how close the candidate text $\bm{y}^{\prime}$ is to the reference text $\bm{y}$. The score measures the precision of matching $n$-grams from the text $\bm{y}^{\prime}$ to $\bm{y}$ by a clipping method to avoid overcounting and adjusting for brevity of $\bm{y}^{\prime}$ if it is shorter than $\bm{y}$. ROUGE-L [41] is the metric for measuring the overlap of sequences of words between the two texts, evaluating how much of the $\bm{y}^{\prime}$ matches with the reference $\bm{y}$ (precision), how much the reference $\bm{y}$ is covered by the candidate $\bm{y}^{\prime}$ (recall), and combine them into an F1 score.
Semantic Similarity: SentenceBert [42] is a variation of the BERT model which is designed to generate high-quality sentence embeddings for the pairs of sentences. We leverage SentenceBert to transform the text $\bm{y}$ and $\bm{y}^{\prime}$ into fixed-length embedding vectors, and calculate the cosine similarity between their embeddings to quantify their semantic similarity.
Probability Measurement: We define two measurements to reflect how the likelihood of generating the original text $\bm{y}$ changes before and after applying the explanation mask. We first define the Probability Ratio (PR) to indicate how less likely to generate the original output $\bm{y}$ when explanatory tokens are masked: $\text{PR}(\bm{x},\bm{y},\bm{m})=\frac{p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})}{p_{\bm{\theta}}(\bm{y}|\bm{x})}$. If the PR score is far below the random baseline, we can conclude that the masked tokens are indeed important to cause the model generating $\bm{y}$. In addition, we also calculate the KL-divergence between these two distributions to measure their difference: $D_{\text{KL}}\big{(}p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})||p_{\bm{\theta}}(\bm{y}|\bm{x})\big{)}$, and a larger score indicates a larger change in generation likelihood and a more accurate explanation.
5.3 Main Results on Faithfulness (Q1)
We evaluate the faithfulness of the explanatory prompt tokens generated by our framework on five aforementioned metrics. Table 1 shows the results of different methods on three datasets by masking $k=3$ identified tokens. For all metrics except the KL-divergence, a lower score is better which is annotated as $\downarrow$. In general, we observe that our method consistently demonstrates better interpretation faithfulness on all datasets compared with baselines, with a clear margin. In particular, we have the following observations demonstrating the advantage of our method:
Table 1: Faithfulness Measurement Results for LLaMA-2(7B-Chat) and Vicuna(7B)
Model Dataset Methods Metric(@3)
[…]
Precision Recall F1
LLaMA-2 (7B-Chat) Alpaca Random 0.601 0.527 0.533 0.522 0.825 0.842 0.134
Captum 0.515 0.409 0.421 0.409 0.680 0.602 0.417
Integrated-Gradient 0.541 0.424 0.435 0.424 0.725 0.726 0.242
XPrompt 0.482 0.388 0.386 0.379 0.642 0.549 0.504
tldr_news Random 0.794 0.742 0.741 0.741 0.923 0.944 0.037
Captum 0.759 0.701 0.703 0.701 0.900 0.910 0.069
Integrated-Gradient 0.713 0.641 0.642 0.640 0.866 0.817 0.149
XPrompt 0.692 0.619 0.610 0.612 0.841 0.604 0.394
MHC Random 0.723 0.617 0.614 0.615 0.787 0.958 0.023
Captum 0.640 0.497 0.493 0.494 0.663 0.760 0.189
Integrated-Gradient 0.646 0.500 0.496 0.498 0.687 0.836 0.117
XPrompt 0.575 0.403 0.405 0.403 0.602 0.710 0.238
Vicuna (7B) Alpaca Random 0.587 0.541 0.552 0.535 0.786 0.891 0.079
Captum 0.466 0.435 0.427 0.418 0.649 0.654 0.347
Integrated-Gradient 0.541 0.495 0.503 0.488 0.744 0.835 0.135
XPrompt 0.433 0.395 0.390 0.377 0.639 0.589 0.459
tldr_news Random 0.781 0.736 0.746 0.737 0.921 0.891 0.091
Captum 0.556 0.470 0.461 0.459 0.775 0.376 0.829
Integrated-Gradient 0.705 0.669 0.672 0.666 0.874 0.821 0.183
XPrompt 0.536 0.456 0.454 0.448 0.772 0.317 1.006
MHC Random 0.715 0.625 0.623 0.623 0.810 0.972 0.012
Captum 0.579 0.438 0.432 0.433 0.627 0.811 0.120
Integrated-Gradient 0.672 0.559 0.555 0.556 0.762 0.941 0.030
XPrompt 0.575 0.431 0.425 0.427 0.620 0.783 0.141
Slight random perturbations on the input prompt do not significantly alter the model’s output, validating the soundness of our approach.
This is evident from the high PR value of around 0.958 and the low KL value of only 0.023 on the MHC dataset when randomly masking tokens. Only when essential tokens are perturbed, particularly when masked, does the generation of content and probability change substantially. The noticeable gaps in these metrics between XPrompt and Random underscore the genuine importance of the masked tokens and their role as counterfactual explanations for the model output.
XPrompt can effectively capture semantically important prompt fragments, and the advantage stands out for long context.
Compared to Captum and Integrated-Gradient, our method generally yields lower values for BLEU and ROUGE-L across all datasets. The discrepancies between XPrompt and Captum are more pronounced for datasets with longer text, such as tldr_news and MHC. Specifically, on the tldr_news dataset, the F1-score of XPrompt is 12.7% lower than Captum for the LLaMA-2 model, and on the MHC dataset, it is 17.85% lower, this all indicates that XPrompt finds and removes the more important token. Furthermore, the SentenceBert results are better on all datasets, indicating larger semantic variations in the generated output after removing tokens by XPrompt.
XPrompt works even better on stronger LLMs.
The gaps between XPrompt and Captum are less apparent on the Vicuna-7B model compared to the LLaMA-2-7B-chat model, which could be attributed to the model’s inherent inference capabilities. Our method is based on the premise that there are textual correlations among the input tokens, allowing the model to infer from the remaining content when a portion of the tokens is masked. If the model has a poor ability to infer the masked token, it may struggle to capture the contextual information. Consequently, the effectiveness of our method is tied to the LLM’s proficiency in inference and understanding.
5.4 Time Efficiency (Q2)
Table 2: Time Efficiency on LLaMA-2(7B-Chat)
Method Dataset
Alpaca tldr_news MHC
Time(s) Captum 1169.648 1727.602 1806.551
XPrompt 15.225 15.397 14.473
[…]
5.5 Qualitative Visualization (Q3)
Figure 2 uses a case study to showcase our method can effectively identify interacted tokens. This figure illustrates the identified explanatory tokens by Captum and our method XPrompt. While there are few overlaps in the found tokens, compared with our method, those found by Captum are mostly not important to the generated response. This demonstrates the Captum’s limited ability to explain the relationship between input prompt and output response, especially considering the textual correlations within input tokens. For instances, the tokens “publishing”, “publish” and “book” have semantic correlations, thus Captum masking one by one individually cannot really eliminate this information complementarily provided by the others, especially when the LLM has a certain context inference ability. Similar issues exist in the second case, where “doctor” and “patient” has semantic interactions. Therefore, treating tokens independently and ignoring their joint influence on the generation is not a favored choice for prompt attribution.
This case study highlights the importance of our formulation of joint attribution, and verifies the effectiveness of our algorithm in discovering important token combinations on the generated output. More examples are shown in Appendix. A.2
Figure 2: Case study for visualizing the explanation of the model responses.
5.6 Ablation Study
Number of Explanatory Tokens.
Figure 3 shows the Probability Ratio (PR) score as the number of masked tokens $k$ changes on each dataset for LLaMA-2(7B-Chat). Our algorithm consistently outperforms other baseline methods even with larger $k$, demonstrating the stable performance and effectiveness of XPrompt. After masking $k=2$ tokens, the value of PR would decrease dramatically, indicating that these two tokens are crucial for generating the output $\bm{y}$ and they act as the triggers that alter the probability distribution of the output. As the number of tokens increase to $k=5$, the PR keeps decreasing but with a less sharp slope. This suggests that the probability of generating the original output $\bm{y}$ is significantly determined by a few predominant tokens.
Figure 3: Probability Ratio with varying number of masked tokens $k$.
Figure 4: Convergence plot of XPrompt.
Number of Sampling Iterations.
We now demonstrate the convergence of our search algorithm. For each sampling iteration, we sample entries in the mask for value swapping; the mask will be updated if the swapping leads to a drop in the generation log-likelihood $\log p_{\bm{\theta}}(\bm{y}|\bm{m}\odot{\bm{x}})$. Figure 4 empirically shows how the log-likelihood decreases as the sampling and mask update continue. The quickly decreasing trend demonstrates that our algorithm is successfully performed to improve the quality of mask in locating the predominant tokens on generating $\bm{y}$.
Effectiveness of Gradient Guidance.
Recall that in our algorithm, we use gradient as guidance to initialize the mask (e.g., $\bm{m}^{(1)}$) and calculate sampling probability (i.e., softmax($\bm{m}^{(n)}\odot\bm{g}$)). To verify the efficiency of gradient guidance, we compare XPrompt with two variants: w/o Initialization that initializes the mask by uniformly random instead of gradient, and w/o Probability that samples swapping entries by uniformly random instead of gradient. Table 4 in the Appendix reports their resulting generation log-likelihood $\log p_{\bm{\theta}}(\bm{y}|\bm{m}\odot{\bm{x}})$ in different iterations.
We observe that without using gradient for initialization, the randomly initialized mask starts from a worse point with a high generation likelihood; and without using gradient to guide the sampling, the mask is updated in less effective direction to explore the search space, resulting in a high generation likelihood in the end. These results show that gradient is a useful and efficient tool for initializing the optimization from a better starting point and guiding the search to a better optimal point.
6 Conclusions
In this study, we introduce XPrompt, an efficient probabilistic search algorithm designed to generate the prompt attributions that elicit the model outputs for the generation tasks. We tackle the challenge of explaining the generation behavior for any given prompt by analyzing the joint effect of the prompt attributions on the output. We frame this explanation task as a discrete optimization problem, which can be efficiently solved by our proposed probabilistic search algorithm. This methodology enables efficient generation of any arbitrary number of explanatory prompt attributions that are deterministic to the generated content. Our framework is rigorously evaluated across extensive language generation tasks, including text summarization, question-answering, and general instruction datasets. The faithfulness of the explanatory prompt attributions is thoroughly analyzed and assessed by comprehensive metrics. The results demonstrate that our proposed method efficiently generates explanatory attributions that faithfully reflect the model’s generation behavior for the specific prompt. These explanatory attributions interact in semantically and jointly influence the output generation. Furthermore, the overall excellent performance of our method on these diverse datasets highlights its remarkable transferability.
Appendix A Appendix
A.1 Dataset Details
We evaluate explanations on three datasets: Alpaca, tldr_news and MHC. Due to the extensive computational cost, we randomly select approaximately 110 data samples with at least 15 words from each datasets. The statistics of the data used in our experiment can be found in Table 3.
  • Alpaca [34] is a dataset with 52000 unique examples consisting of instructions and demonstrations generated by OpenAI’s text-davinci-003 engine. Each example in the dataset includes an instruction that describes the task the model should perform, accompanied by optional input context for that task. In our experiment, we randomly select a subset of these examples for verification purposes.
  • tldr_news [35] dataset is constructed by collecting a daily tech newsletter. For every piece of data, there is a headline and corresponding content extracted. The task is to ask the model to simplify the extracted content and then generate a headline from the input.
  • mental_health_counseling (MHC) [36] includes broad pairs of questions and answers derived from online counseling and therapy platforms. It covers a wide range of mental-health related questions and concerns, as well as the advice provided by the psychologists. It is utilized for used to fine-tuning the model to generate the metal health advice. Here, we prompt the model to generate an advice based on the provided question.
Dataset Length Number Prompt Examples
Alpaca 16-137 108 Pretend you are a project manager of a construction company. Describe a time when you had to make a difficult decision.
tldr_news 14-138 120 Reddit aired a five-second long ad during the Super Bowl. The ad consisted of a long text message that hinted at the GameStop stocks saga. A screenshot of the ad is available in the article.
MHC 18-478 100 I cannot help myself from thinking about smoking. What can I do to get rid of this addiction?
Table 3: Data Statistics
A.2 Case Study
In this section, we visualize three examples which are selected from the dataset Alpaca, tldr_news, and MHC respectively, shown in Figure. 5. The tokens in the red frame are the $k$ explanatory tokens extracted by the model LLaMA-2(7B-Chat), and the output responses framed in blue are the most informative results selected by human. In the first data sample, the recognized token: "brains" and "cognitive", has some semantic relationship intuitively. It is a common sense the cognitive ability is underpinned by the brain’s function and structure. There are numerous reports and studies focused on the functionally relationship between the brain and coginitive ability [43]. Due to the extensive amount of data on which the LLMs are trained, the model can integrate their relationship which is reflected in its response.
As for the second example, the visualized tokens are almost related to the "book" and "publish", which aligns with the theme of the output content. Removing either "publishing" or "publish" would not greatly alter the meaning of the input prompt. This example also shows that our method could detect some fix expressions like "whether…or..", whose relation is recognized by the token "vs." in the output.
In the last example, we could figure out the model infers or "guesses" that the kid feels frustrated and helpless, a sentiment partly related to the token "never". The generated output does not mention the reason to explain how the model makes such inference. From this point, our method could generate human-intelligible explanations to assist the users in identifying which parts of the prompts fundamentally lead to the undesired response, providing guidance on how to revise the prompts effectively.
Figure 5: Case study for visualizing the top-$k$ tokens in each example.
Figure 6: Case study for visualizing the variation of the model responses after the specific tokens are masked.
In Figure 6, we also illustrate the changes in the model response as the targeted top-$k$ token is removed from the input prompt. Compared with the top-$k$ tokens identified by our method, those found by Captum are mostly not important to the response. This demonstrates the Captum’s limited ability to explain the relationship between input prompt and output response with the consideration of the textual information. Moreover, when masking the token "publishing" or "publish", as identified by our method XPrompt, the generated outputs do not exhibit significant variations in either word frequency or semantics. However, substantial changes in the generated outputs occur only when both tokens are simultaneously masked. This highlights the effectiveness of our methodology in explaining the joint effect of the prompt attributions on the generated output.
A.3 Gradient Design
The Table. 4 displays the log-likelihood without using the gradient as the guidance to initialize the optimization process or to do sampling at different stage when searching for the optimal results.
Methods Number of Iterations
1 5 10 15 20 30 40 50
XPrompt -42.448 -51.661 -56.378 -59.852 -62.303 -64.812 -67.274 -69.426
OO w/o Initialization -33.906 -44.555 -50.828 -55.652 -58.483 -62.694 -64.793 -66.415
OO w/o Probability -41.759 -49.867 -53.784 -57.107 -59.639 -63.339 -65.848 -67.496
Table 4: Ablation study showing the log-likelihood without using gradient for initialization or sampling with different number of iterations.
A.4 Limitations
One possible limitation of our work is that while the explanations demonstrate the counterfactual correlation between the input prompt and the output sequence, the underlying generation mechanism remains a black box and could not be explained by our method, which we believe is an important and challenging problem. In the future work, we aim to extend the proposed framework to explore more aspects of the generation mechanism in LLMs.