### v2

# JoPA: Explaining Large Language Model’s Generation via Joint Prompt Attribution

Abstract
Large Language Models (LLMs) have demonstrated impressive performances in complex text generation tasks. However, the contribution of the input prompt to the generated content still remains obscure to humans, underscoring the necessity of understanding the causality between input and output pairs. Existing works for providing prompt-specific explanation often confine model output to be classification or next-word prediction. Few initial attempts aiming to explain the entire language generation often treat input prompt texts independently, ignoring their combinatorial effects on the follow-up generation. In this study, we introduce a counterfactual explanation framework based on Joint Prompt Attribution, JoPA, which aims to explain how a few prompt texts collaboratively influences the LLM’s complete generation. Particularly, we formulate the task of prompt attribution for generation interpretation as a combinatorial optimization problem, and introduce a probabilistic algorithm to search for the casual input combination in the discrete space. We define and utilize multiple metrics to evaluate the produced explanations, demonstrating both the faithfulness and efficiency of our framework. Our code is available at  https://github.com/yuruic/JoPA.
1 Introduction
Large Language Models (LLMs), such as GPT4 (achiam2023gpt, ), LLaMA (touvron2023llama, ) and Claude (anthropic_2024_claude, ), have shown excellent performance in various natural language generation tasks including question answering, document summarization, and many more. Despite the great success of LLMs, we still have very limited understanding of the LLM generation behavior – which parts in the input cause the model to generate a certain sequence. Unable to explain the causality between the input prompt and the output generation could cause failure in recognizing potential unintended consequences, such as harmful response (dan, ; autodan-liu, ; gcg, ; autodan-zhu, ) and biased generation (wang2023decodingtrust, ) attributed to a specific malicious description in the input. These issues undermine human trust in model usage, thus highlighting a pressing need for developing an interpretation tool that attributes how an input prompt leads to the generated content.
Explaining LLM generation through prompt attribution involves extracting the most influential prompt texts on the model’s entire generation procedure, a realm that remains relatively under-explored in current research. While extensive works on input attribution are proposed for text classification interpretation (chen2020learning, ; modarressi2023decompx, ) and next-word generation rationale (zhao2024reagent, ; vafa2021rationales, ), they can not be directly applied to explain the full generation sequence due to its complicated joint probability landscape and the autoregressive generation procedure. The complexity for interpreting LLM generation compounds as the model size increases. Another line of works involves prompting LLMs to self-explain their behaviors (wei2022chain, ). This method relies on the model’s innate reasoning capabilities, although current findings suggest that these capabilities may not always be faithful (Turpin2023LanguageMD, ; xu2024hallucination, ).
There are limited existing attempts focusing on explaining the relationship between the input prompts and the complete generated sequence. The most relevant approach, Captum (captum, ), sequentially determines the importance score for each token by calculating the variations in the joint probability of generating the targeted output sequence when the token is dropped from the model input. While being straightforward, this approach treats tokens as independent features, ignoring their joint semantic influence on the generated output. In fact, tokens may contain overlapping or complementary information. For example, given the input prompt: “Write a story about the doctor and his patient”, the most influential components are “doctor” and “patient”. Individually removing either of these words would not significantly alter the generated output, resulting in inaccurately low importance score for each of them. This is caused by the semantic interaction among these components, allowing the model to infer the meaning of the omitted one. Existing attribution methods that ablate each token individually fail to capture such combinatorial effect. A straightforward remedy might involve exhaustively assessing all possible combinations to observe the variations in the model generations, which however is impractical due to the vast search space with long-context input prompts.
To efficiently search the space for generating accurate prompt explanations, we develop our framework, Joint Prompt Attribution (JoPA), which provides a counterfactual explanation to highlight which components of input prompts have the fundamental effect on the generated context by solving a combinatorial optimization problem. We aim to explain the generation behavior of model outputs for any given prompt while taking the joint effects of the prompt components into account. Assuming that removing the essential parts of the prompt would result in a significant variation in the model’s output, we propose the novel objective function and formulate our task of providing faithful counterfactual explanations for the input prompt as an optimization problem. To quantify the influence of token combinations on the prompt on generations, we incorporate a mask approach for joint prompt attribution. Thus, our goal of extracting the explanations has been converted to finding the optimal mask of the input prompt. We solve this problem by a probabilistic search algorithm, equipped with gradient guidance and probabilistic updates for efficient exploration in the discrete solution space. Our main contributions can be summarized as follows:
  • We propose a general interpretation scheme for LLM generation task that attributes input prompts to the entire generation sequence, considering the joint influence of input token combinations. This motivation naturally formulates a combinatorial optimization problem to explain generation with the most influential prompt texts.
  • We demonstrate JoPA, an efficient probabilistic search algorithm to solve the optimization problem. JoPA works by searching better token combinations that lead to larger generation changes. It takes the advantage of both the gradient information and the probabilistic search-space strategy, thereby achieving an efficient prompt interpretation tool.
  • Our framework demonstrates strong performance on language generation tasks including text summarization, question-answering, and general instruction datasets. The faithfulness of our explanations is assessed using metrics on generation probability, word frequency, and semantic similarity, confirming the methods’ transferability and effectiveness across tasks. Additionally, the explanations highlight the framework’s potential to enhance model safety and efficiency.
2 Related Work
While there are extensive works devoted to explaining language models in the context of text classification tasks (shi2022nearest, ; han2021ptr, ; shi2023chatgraph, ), relatively few attempts (zhao2024reagent, ; vafa2021rationales, ) are proposed to investigate the importance of prompt texts on the entire generation procedure especially for LLMs. This demonstrates a research gap that this work aims to fill in.
2.1 Explaining Language Generation
There are many work explaining the predictions generated by LLMs by measuring the importance of the input features to the model’s prediction on the classification tasks. One group of studies perturb the specific input by removing, masking, or altering the input features, and evaluate the model prediction changes (10.1145/3461702.3462597, ; wu-etal-2020-perturbed, ). The other group of works, such as integrated gradients (IG) (sundararajan2017axiomatic, ), first-derivative saliency (li-etal-2016-visualizing, ), and mixed partial derivatives (tsang2020does, ) leverage the gradients of the output with respect to the input to determine the input feature importance. Although Archipelago (tsang2020does, ) explains the feature attributions by considering the combined effects of the input attributions, it targets the multilabel classification task and relies on the neutral baseline. As for the generation tasks, Diffmask learns the differentiable mask for each layer of the BERT model, but it aims to analyze how decisions are formed across network hidden layers by a simple probing classifier. In contrast, our framework targets on offering insights into the relationship between the input prompt and model generations by employing the mask to highlight the essential prompt attributions.
Moreover, a few studies utilize the surrogate model to explain the individual predictions of the black-box models, and the representative method is called LIME (ribeiro2016should, ). As for explaining the model’s generation behavior, Captum (captum, ) calculates the token’s importance score by sequentially measuring the contribution of input token to the output, which lacks accounting for the semantic relationships between tokens. Another work, ReAGent (zhao2024reagent, ), focuses on the next-word generation task, computing the importance distribution for the next token position. This method ignores the contextual dependencies in the generated output and could not adequately account for dynamically changing generations. Our framework aims to interpret the joint effects of the input prompts on the entire output contexts with considering of the textual information covered the input prompt.
2.2 Self-explaining by Prompting
As language models increase in scale, prompting-based models demonstrate remarkable abilities in reasoning (brown2020language, ), creativity (oppenlaender2022creativity, ), and adaptability across a range of tasks (khattak2023maple, ). However, the complex reasoning processes of these models remain elusive and require tailored paradigm to better understand the prompting mechanism. For instance, the chain-of-thought (CoT) paradigm could explain the LLM behaviors by prompting the model to generate the reasoning chain along with the answers (wei2022chain, ), as pre-trained LLMs have demonstrated a certain ability to self-explain their behaviors. However, recent studies have also suggested that the reasoning chain does not guarantee faithful explanations of the model’s behavior (jacovi-goldberg-2020-towards, ) and the final answer might not always follow the generated reasoning chain (Turpin2023LanguageMD, ). xLLM (chuang2024large, ) improves explanation fidelity from LLMs using a fidelity evaluator, tailored for classification tasks. We focus on explaining LLM generation by analyzing input prompt attributions to output content, independent of the model’s suboptimal reasoning abilities.
3 Preliminaries
We first introduce necessary notions for the LLM generation process, and then discuss limitations of prior attempts explaining the entire generation via prompt attribution.
LLM Generation Notions
Denote a specific input prompt as a sequence of tokens $\bm{x}=(x_{1},\dots,x_{T})$, $x_{i}\in\{1,2,\dots,|V|\}$, where $|V|$ represents the vocabulary size, $T$ is the length of the input sequence and the set of all token indices is $\mathcal{I}=\{1,2,\dots,T\}$. The corresponding generated output $\bm{y}$ could be represented as a sequence of tokens $\bm{y}=(y_{1},\dots,y_{S})$ with $y_{j}\in\{1,2,\dots,|V|\}$. The output tokens $\{y_{i}\}_{i=1}^{S}$ are generated from the LLM $f_{\bm{\theta}}$ parameterized by $\bm{\theta}$ in an autoregressive manner with the probability $p_{\bm{\theta}}(\bm{y}|\bm{x})$ as:
$p_{\bm{\theta}}(\bm{y}|\bm{x})=p_{\bm{\theta}}(y_{1}|\bm{x})\prod_{i=1}^{S-1}p_{\bm{\theta}}(y_{i+1}|\bm{x},y_{i}).$ (1)
The probability of generating the output text $\bm{y}$ given the input prompt $\bm{x}$ illustrates that the coherence and generation of the output texts heavily rely on the input prompt $\bm{x}$, indicating an implicit causal relationships between the input prompt $\bm{x}$ and the output $\bm{y}$.
Limitation of Prior Interpretation Attempt
There are limited prior works about explaining the relationship between the individual input prompts and the entire generated sequence, and their faithfulness is limited by treating input tokens independently. Specifically, Captum (captum, ), calculates the importance score for each token by slightly perturbing the input $\bm{x}$ at the $i$-th token:
$G_{i}(\bm{x};\bm{\theta})=p_{\bm{\theta}}(\bm{y}|\bm{x})-p_{\bm{\theta}}(\bm{y}|\bm{x}_{\mathcal{I}\setminus\{i\}}),$ (2)
where $\mathcal{I}$ denotes all token indices. Through Eq. (2), one can calculate an attribution score for each token $i$ indicating its importance towards the original generation result $\bm{y}$. Such method is direct and simple, however, it treats tokens as independent features, ignoring their semantic interaction and joint effect on the generated output. To make it more obvious, let’s recall the previous example of “Write a story about a doctor and his patient.”. Note that since “doctor” and “patient” are semantically correlated, and removing any of the two words would not have a significant influence on the generated response, their corresponding importance score will not be too high. On the contrary, the most important token detected by this approach would be “his”, which is clearly not ideal.
These observations motivate a new prompt attribution method that accounts for semantic correlations among tokens. We assume LLM outputs are primarily influenced by a subset of tokens acting jointly, while others serve as auxiliary information. These subset of tokens, which shape the output collectively, are identified as explanatory tokens, contributing to responses through joint effects rather than independently.
4 Proposed Method
In this section, we propose Joint Prompt Attribution (JoPA), a simple yet effective framework designed for generative tasks to explain the attribution of the input prompt by solving a discrete optimization problem. We start with formulating the general objective for the discrete optimization problem, and then we introduce our proposed probabilistic search algorithm for solving the problem.
Figure 1: Overview of JoPA. Left: Demonstrating the pipeline of the algorithm. Right: Illustrating the process of mask $\bm{m}$ sampling.
Problem Formulation for Joint Prompt Attribution
In order to explain language generation via prompt attribution, instead of treating the prompt tokens independently as in previous works (captum, ), we propose to evaluate the joint effect of $k$ prompt tokens on the generated sequence.
Specifically, consider a binary prompt mask $\bm{m}=(m_{1},\dots,m_{T})$ in which $m_{i}\in\{0,1\}$ indicates whether the $i$-th token is important or not. The joint probability of generating the original output sequence $\bm{y}$ given an input masked context $\bm{m}\odot\bm{x}$ is computed as $p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})$, where $\odot$ denotes the Hadamard product. In our paper, we aim to identify a binary mask $\bm{m}$, which is a (discrete) learnable parameter, that maximizes the discrepancy in the probability of generating the same output $\bm{y}$ when comparing the masked input $\bm{m}\odot\bm{x}$ to the original input $\bm{x}$. A large discrepancy indicates that the masked token combination are the most influential components for generating $\bm{y}$, thus should jointly serve as the attributed explanation. Consequently, this involves training the binary mask $\bm{m}$ to optimize the following objective function:
$\max_{\bm{m}\in\{0,1\}^{T}}$ $\mathcal{L}(\bm{m},\bm{x},\bm{y};\bm{\theta})$ (3)
$=p_{\bm{\theta}}\left(\bm{y}|\bm{x}\right)-p_{\bm{\theta}}\left(\bm{y}|\bm{m}\odot\bm{x}\right)$
$\text{s.t.}\quad|\bm{m}|_{1}$ $=T-k,$
where $k$ represents the number of explanatory tokens. Intuitively, optimizing the objective outlined in Eq. (3) suggests that we want to find the top $k$ important tokens which, if they are masked, lead to a substantial variation in model’s output probabilities. Compared with prior methods for individual token attribution stated in Eq. (2), our formulation in Eq. (3) measures the joint attribution of a subset of tokens, capturing token interactions to enable more accurate prompt explanations.
Challenges in Solving Eq. (3)
Note that Eq. (3) is a constraint discrete optimization problem that is non-trivial to solve. One naive solution would be transforming the discrete optimization problem into a continuous one, i.e., let $\bm{m}\in[0,1]^{T}$, and formulate the constraint into a regularization term. Then one can adopt traditional gradient descent based optimization solutions to solve the problem. However, such a strategy would require extensive gradient calculations and backward steps on the original inputs, which can be inefficient in practice especially for LLMs. Moreover, the obtained continuous mask is not the final explanation we want. In fact, the approximation error when transforming the continuous mask back into the discrete space can also be quite significant, leading to worse performances. Another straightforward solution involves searching through all possible token combinations exhaustively. However, this could be impractical as well due to the enormous search space especially while facing long-context inputs. Therefore, we hope to develop a new method that adopts a search-based strategy to simplify the algorithm design and satisfy our constraints, while also leveraging gradient information for efficient optimization.
Algorithm 1 Explainable Prompt Generator: JoPA
Input: Input tokens $\bm{x}$, output tokens $\bm{y}$, the integer $k$ denoting the number of explanatory tokens, $1\leq k\leq T$, input mask $\bm{m}^{(0)}=\textbf{1}$, and the sampling numbers $N$.
Output: Optimal mask $\bm{m}^{(N)}$
[…]
2:Set $\bm{m}^{(1)}$ as the $\text{top-}k$ value mask for $\bm{g}$: $\bm{m}^{(1)}=\bm{m}^{(0)};m_{i}^{(1)}=0,\forall i\in\text{argtop-}k(\bm{g})$
3:for $n=1$ to $N$ do
4:  Sample $l\sim\text{softmax}\big{(}\bm{m}^{(n)}\odot\bm{g}\big{)}$
5:  Sample $v\sim\text{softmax}\big{(}(\mathbf{1}-\bm{m}^{(n)})\odot\bm{g}\big{)}$
6:  $\bm{m}^{\text{tmp}}=\bm{m}^{(n)}.\text{copy}()$; switch the value of ${m}^{\text{tmp}}_{l}$ and ${m}^{\text{tmp}}_{v}$
7:  if $p_{\bm{\theta}}(\bm{y}|{\bm{m}^{\text{tmp}}}\odot\bm{x})<p_{\bm{\theta}}(\bm{y}|{\bm{m}^{(n)}}\odot\bm{x})$ then
8:   $\bm{m}^{(n+1)}=\bm{m}^{\text{tmp}}$
9:  else
10:   $\bm{m}^{(n+1)}=\bm{m}^{(n)}$
11:  end if
12:end for
Proposed Probabilistic Search Algorithm
To tackle this challenge, we propose JoPA, a novel explainable prompt generator, for efficiently obtaining an optimal solution for Eq. (3). We summarize our JoPA in Algorithm 1. The high-level pipeline is illustrated in Figure 1 left: we initialize and maintain exactly $k$ entries in the mask $\bm{m}$ to be zero to enforce the constraint and capture their joint influence of being masked; the mask is then iteratively updated by sampling indexes for value swapping, searching towards the direction with increased generation loss $\mathcal{L}(\bm{m},\bm{x},\bm{y};\bm{\theta})$. At the core of this pipeline is the sampling and update of the discrete mask, which demands an efficient exploration in the vast search space. Figure 1 right shows this step, which is featured by the following two essential components, the gradient-guided masking and the probabilistic search update.
Gradient-Guided Masking: Trivial solutions that set the mask $\bm{m}$ by uniformly random could cost massive sampling to hit the right optimization direction. Gradient is a common indicator of feature importance, as evidenced in existing practices (ebrahimi2018hotflip, ; gcg, ; shin-etal-2020-autoprompt, ). Therefore, we propose to use gradient as a guidance to set and sample the mask for more efficient optimization. Specifically, we begin with the binary mask $\bm{m}^{(0)}=\bm{1}$ which indicates that all tokens in the input $\bm{x}$ are marked as non-explanatory ones. Then we compute the gradient $\nabla_{\bm{m}^{(0)}}\mathcal{L}(\bm{m}^{(0)},\bm{x},\bm{y};\bm{\theta})$ of the loss function in Eq. (3), and denote the magnitude of gradients as $\bm{g}=|\nabla_{\bm{m}^{(0)}}\mathcal{L}(\bm{m}^{(0)},\bm{x},\bm{y};\bm{\theta})|$. Note that the components with larger gradient magnitudes in $\bm{g}$ imply that altering the corresponding tokens could result in a sharper change to the generated outputs. In order to initialize the explanatory $k$ tokens guided by the gradient magnitude, we set the binary mask $\bm{m}^{(1)}$ where $m^{(1)}_{i}=0$ indicates $g_{i}$ is the top-$k$ value in $\bm{g}$. The gradient guidance and mask initialization are obtained following Line 1-2 in the algorithm.
[…]
JoPA combines gradient information with a search-space strategy, achieving greater efficiency than using either alone. It requires only one gradient calculation step and avoids converting between discrete and continuous masks. Gradients offer an efficient search direction and reliable sampling probabilities. Theoretical insights are detailed in Appendix A.1.
5 Experiment
This section aims to verify the effectiveness and efficiency of our proposed framework for interpreting the LLMs on the generation task. We conduct the experiments to answer the following questions:
  • Q1: Do the generated prompt attributions play a predominant role in the model’s generation, thereby serving as faithful counterfactual explanations for the generated output?
  • Q2: Is our interpretation algorithm efficient for practical usage?
  • Q3: Could the proposed algorithm effectively identify a combination of tokens that impose joint influence on the model generation?
We provide quantitative studies to evaluate the faithfulness of the explanatory prompt fragments generated by JoPA, comparing with existing interpretation baselines and ablation variants.
5.1 Experiment Settings
Models & Baselines
In the experiment, we employ two LLMs as our targeted $f_{\bm{\theta}}(\cdot)$: LlaMA-2 (7B-Chat) (touvron2023llama, ) and Vicuna (7B) (Zheng2023JudgingLW, ). There are relatively few methods attributing prompts on the entire language generation, and we choose random removal (Random), Integrated-Gradient (sundararajan2017axiomatic, ), averaged attentions across all layers (pruthi2019learning, )(Attention), last-layer attention (zhao2024reagent, )(Last-Attention) and Captum (captum, ) as our baseline for comparison. We also compare JoPA with ReAGent (zhao2024reagent, ), with the results presented in Appendix A.13. We implement these models using the PyTorch framework and pretrained weights from the transformers Python library (wolf-etal-2020-transformers, ), and conduct our experiments on an Nvidia RTX A6000-48GB platform with CUDA version 12.0.
Datasets
We employ three distinct text generation datasets: Alpaca (alpaca, ), tldr_news (tldr_news, ), and MHC (amod_2024, ), to evaluate the effectiveness of our method across various generation tasks. To capture the joint influence of prompts on model outputs, we use long-context prompts (15+ words) rather than simple sentences, as they convey more information and exhibit higher textual correlation. For evaluation, 110 samples are randomly selected from each publicly available dataset, detailed in AppendixA.2
5.2 Evaluation Metrics
Faithfulness scores are a key metric for assessing the quality of explanations, with faithful explanations accurately reflecting the model’s decision-making process (jacovi-goldberg-2020-towards, ). Studies (7552539, ; hooker2019benchmark, ) suggest that if a certain input tokens are truly important, their removal should lead to a more significant change in model output than the removal of random tokens (madsen-etal-2022-evaluating, ). Therefore, after removing explanatory tokens identified by a faithful method, the model’s new generation would significantly differ from using the original output. Moreover, the input prompt with masking these explanatory tokens are less likely to reproduce the original model response $\bm{y}$, indicating a lower value of $p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})$.
To quantitatively evaluate explanation faithfulness, we measure the change of model generation behavior from two perspectives. First, we compare the original generation $\bm{y}=f_{\bm{\theta}}(\bm{x})$ with the new generation $\bm{y}^{\prime}=f_{\bm{\theta}}(\bm{m}\odot\bm{x})$, measuring similarity in word frequency and semantics. Lower similarity indicates greater change, suggesting better explanations. Second, we evaluate the likelihood of the original output $\bm{y}$ under the masked prompt, $p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})$. Lower likelihood reflects more effective masks that prevent the model from replicating the original output. Detailed metric definitions are explained below.
Word Frequency: BLEU (bleu, ) evaluates how closely the candidate text $\bm{y}^{\prime}$ matches the reference $\bm{y}$ by measuring n-gram precision with clipping to avoid overcounting and brevity adjustments for shorter candidates. ROUGE-L (rouge, ) measures sequence overlap between $\bm{y}^{\prime}$ and $\bm{y}$, assessing precision, recall, and their F1 score.
Semantic Similarity: SentenceBert (reimers2019sentencebert, ) is a variation of the BERT model which is designed to generate high-quality sentence embeddings for the pairs of sentences. We leverage SentenceBert to transform the text $\bm{y}$ and $\bm{y}^{\prime}$ into fixed-length embedding vectors, and calculate the cosine similarity between their embeddings to quantify their semantic similarity.
Probability Measurement: We define two measurements to reflect how the likelihood of generating the original text $\bm{y}$ changes before and after applying the explanation mask. We first define the Probability Ratio (PR) to indicate how less likely to generate the original output $\bm{y}$ when explanatory tokens are masked: $\text{PR}(\bm{x},\bm{y},\bm{m})=\frac{\tilde{p}_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})}{\tilde{p}_{\bm{\theta}}(\bm{y}|\bm{x})}$, where $\tilde{p}(\bm{y}|\bm{x})=p(\bm{y}|\bm{x})^{1/s}$ is the length-normalized generation probability. If the PR score is far below the random baseline, we can conclude that the masked tokens are indeed important to cause the model generating $\bm{y}$. In addition, we also calculate the KL-divergence between these two distributions to measure their difference: $D_{\text{KL}}\large(p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})||p_{\bm{\theta}}(\bm{y}|\bm{x})\large)$, and a larger score indicates a larger change in generation likelihood and a more accurate explanation.
5.3 Main Results on Faithfulness (Q1)
We evaluate the faithfulness of explanatory prompt tokens using five metrics. Table 1 presents results on three datasets by masking $k=3$ identified tokens, with the larger dataset results in Appendix A.8. For all metrics except KL-divergence ($\downarrow$ indicates lower is better), our method consistently outperforms baselines across all datasets with a clear margin, demonstrating superior interpretative faithfulness. Key observations highlight our method’s advantages:
Model Dataset Methods Metric(@3)
[…]
Precision Recall F1
LlaMA-2 (7B-Chat) Alpaca Random 0.601 0.527 0.533 0.522 0.825 0.842 0.134
Last-Attention 0.533 0.423 0.452 0.432 0.672 0.770 0.202
Attention 0.547 0.447 0.460 0.448 0.721 0.791 0.170
Captum 0.515 0.409 0.421 0.409 0.680 0.602 0.417
Integrated-Gradient 0.541 0.424 0.435 0.424 0.725 0.726 0.242
JoPA 0.484 0.388 0.386 0.379 0.642 0.549 0.504
tldr_news Random 0.794 0.742 0.741 0.741 0.923 0.944 0.037
Last-Attention 0.747 0.683 0.685 0.683 0.876 0.869 0.108
Attention 0.767 0.703 0.710 0.706 0.900 0.899 0.077
Captum 0.759 0.701 0.703 0.701 0.900 0.910 0.069
Integrated-Gradient 0.713 0.641 0.642 0.640 0.866 0.817 0.149
JoPA 0.692 0.619 0.610 0.612 0.841 0.604 0.394
MHC Random 0.723 0.617 0.614 0.615 0.787 0.958 0.023
Last-Attention 0.660 0.529 0.525 0.526 0.736 0.907 0.023
Attention 0.665 0.536 0.538 0.531 0.745 0.916 0.056
Captum 0.640 0.497 0.493 0.494 0.663 0.760 0.189
Integrated-Gradient 0.646 0.500 0.496 0.498 0.687 0.836 0.117
JoPA 0.575 0.403 0.405 0.403 0.602 0.701 0.246
Vicuna (7B) Alpaca Random 0.587 0.541 0.552 0.535 0.786 0.891 0.079
Last-Attention 0.475 0.423 0.436 0.415 0.672 0.840 0.113
Attention 0.486 0.447 0.469 0.441 0.683 0.831 0.140
Captum 0.466 0.435 0.427 0.418 0.649 0.654 0.347
Integrated-Gradient 0.541 0.495 0.503 0.488 0.744 0.835 0.135
JoPA 0.433 0.395 0.390 0.377 0.639 0.589 0.459
tldr_news Random 0.781 0.736 0.746 0.737 0.921 0.891 0.091
Last-Attention 0.746 0.707 0.718 0.708 0.898 0.876 0.114
Attention 0.621 0.514 0.515 0.509 0.817 0.714 0.294
Captum 0.556 0.470 0.461 0.459 0.775 0.376 0.829
Integrated-Gradient 0.705 0.669 0.672 0.666 0.874 0.821 0.183
JoPA 0.536 0.456 0.454 0.448 0.772 0.317 1.006
MHC Random 0.715 0.625 0.623 0.623 0.810 0.972 0.012
Last-Attention 0.684 0.570 0.573 0.570 0.773 0.947 0.028
Attention 0.685 0.581 0.581 0.580 0.775 0.950 0.026
Captum 0.579 0.438 0.432 0.433 0.627 0.811 0.120
Integrated-Gradient 0.672 0.559 0.555 0.556 0.762 0.941 0.030
JoPA 0.575 0.431 0.425 0.427 0.620 0.783 0.141
Table 1: Faithfulness Measurement Results for LlaMA-2 (7B-Chat) and Vicuna (7B)
Slight random perturbations on the input prompt do not significantly alter the model’s output, validating the soundness of our approach.
This is supported by a high PR value of around 0.958 and low KL value of only 0.023 on the MHC dataset when masking random tokens. Substantial changes occur only when essential tokens are masked, highlighting their importance as counterfactual explanations. The metric gaps between JoPA and Random emphasize this, with performance stability detailed in Appendix A.9.
JoPA can effectively capture semantically important prompt fragments, and the advantage stands out for long context.
Compared to Captum and Integrated-Gradient, JoPA achieves lower BLEU and ROUGE-L values across all datasets, especially for longer texts like tldr_news and MHC. On tldr_news, the F1-score of JoPA is 12.7% lower, and on MHC, 17.85% lower than Captum for the LlaMA-2 (7B-Chat) model, indicating it identifies and removes more important tokens. Additionally, higher SentenceBert scores across datasets suggest greater semantic variation in outputs after token removal by JoPA.
JoPA works even better on stronger LLMs.
The gaps between JoPA and Captum are less apparent on the Vicuna (7B) model compared to the LlaMA-2 (7B-Chat) model, which could be attributed to the model’s inherent inference capabilities. JoPA is based on the premise that there are textual correlations among the input tokens, enabling inference from remaining content when tokens are masked. Models with weaker inference abilities may struggle to capture context, making our method’s effectiveness dependent on the LLM’s proficiency in inference and understanding.
5.4 Time Efficiency (Q2)
Method Dataset
Alpaca tldr_news MHC
Time(s) Captum 1169.648 1727.602 1806.551
JoPA 15.225 15.397 14.473
Table 2: Time Efficiency on LLaMA-2 (7B-Chat)
In Table 2, we compare the average time cost of JoPA and Captum for generating $k=3$ explanatory tokens for each prompt instance. Note that since Captum’s design requires sequentially appending the next token to the input prompt to re-generate the new output, the time consumption of Captum would increase significantly when the prompt is long. As the average length of MHC is longer than the Alpaca shown in Table 3, the computational time increases from 1169.648ss to 1806.551ss for Captum, which is quite inefficient and impractical. Our algorithm efficiently solves the combinatorial optimization problem using a probabilistic search, significantly reducing computational costs for explanation generation. Since it requires only a certain number of gradient-guided search steps, computation time remains consistent regardless of prompt length, as shown in Table 3.
Figure 2: Case study for visualizing the explanation of the model responses.
5.5 Qualitative Visualization (Q3)
Figure 2 uses a case study to showcase our method can effectively identify interacted tokens. This figure illustrates the identified explanatory tokens by Captum and our method JoPA. While there are few overlaps in the found tokens, compared with our method, those found by Captum are mostly not important to the generated response. This demonstrates the Captum’s limited ability to explain the relationship between input prompt and output response, especially considering the textual correlations within input tokens. For instances, the tokens “publishing”, “publish” and “book” have semantic correlations, thus Captum masking one by one individually cannot really eliminate this information complementarily provided by the others, especially when the LLM has a certain context inference ability. Similar issues exist in the second case, where “doctor” and “patient” has semantic interactions. Therefore, treating tokens independently and ignoring their joint influence on the generation is not a favored choice for prompt attribution. This case study underscores the importance of joint attribution and demonstrates our algorithm’s effectiveness in identifying key token combinations. Additional examples are in Appendix A.3. The algorithm also shows promise for enhancing model safety and aiding diagnosis, detailed in Appendix A.4.
6 Conclusions
We introduce JoPA, a probabilistic search algorithm that efficiently identifies prompt attributions explaining model outputs in generation tasks. By framing the explanation as a discrete optimization problem, JoPA generates faithful and deterministic attributions that influence outputs. Rigorous evaluation on tasks like summarization, question-answering, and instruction datasets shows JoPA’s effectiveness, faithfulness, and transferability.
7 Limitations
One possible limitation of our work is that while the explanations demonstrate the counterfactual correlation between the input prompt and the output sequence, the underlying generation mechanism remains a black box and could not be explained by our method, which we believe is an important and challenging problem. In the future work, we aim to extend the proposed framework to explore more aspects of the generation mechanism in LLMs.
Appendix A Appendix
A.1 Theoretical Insight
Here we could prove that our algorithm can theoretically converge to the local optima given enough iterations. Define that a solution $\bm{m}^{*}\in\{0,1\}^{T}$ is the local optima, if we have $p_{\bm{\theta}}(y|\bm{m}^{*}\odot\bm{x})\leq p_{\bm{\theta}}(y|\bm{m}\odot\bm{x})$ for all $\bm{m}$ in the one-swap neighborhood of $\bm{m}^{*}$, namely $\bm{m}^{*}-\bm{m}||_{0}=2$ (meaning $\bm{m}$ differs from $\bm{m}^{*}$ by one swap, as they are constrained to have $k$ zero entries). It could be proved that the output of the algorithm $\hat{\bm{m}}$ is the local optima by contradiction that with sufficient iterations. Suppose $\hat{\bm{m}}$ is not the local optima, there must be a point $\bm{m}^{\prime}$ in its neighbor satisfying $\bm{m}^{\prime}-\hat{\bm{m}}||_{0}=2$, such that $p_{\bm{\theta}}(y|\bm{m}^{\prime}\odot x)<p_{\bm{\theta}}(y|\hat{\bm{m}}\odot\bm{x})$. This means our algorithm can still find a better solution by sampling another swap in further iterations and thus $\hat{\bm{m}}$ is not the output of our algorithm, leading to a contradiction to the assumption.
A.2 Dataset Details
We evaluate explanations on three datasets: Alpaca, tldr_news and MHC. Due to the extensive computational cost, we randomly select approximately 110 data samples with at least 15 words from each datasets. The statistics of the data used in our experiment can be found in Table 3.
  • Alpaca alpaca is a dataset with 52000 unique examples consisting of instructions and demonstrations generated by OpenAI’s text-davinci-003 engine. Each example in the dataset includes an instruction that describes the task the model should perform, accompanied by optional input context for that task. In our experiment, we randomly select a subset of these examples for verification purposes.
  • tldr_news tldr_news dataset is constructed by collecting a daily tech newsletter. For every piece of data, there is a headline and corresponding content extracted. The task is to ask the model to simplify the extracted content and then generate a headline from the input.
  • mental_health_counseling (MHC) amod_2024 includes broad pairs of questions and answers derived from online counseling and therapy platforms. It covers a wide range of mental-health related questions and concerns, as well as the advice provided by the psychologists. It is utilized for used to fine-tuning the model to generate the metal health advice. Here, we prompt the model to generate an advice based on the provided question.
Dataset Length Number Prompt Examples
Alpaca 16-137 108 Pretend you are a project manager of a construction company. Describe a time when you had to make a difficult decision.
tldr_news 14-138 120 Reddit aired a five-second long ad during the Super Bowl. The ad consisted of a long text message that hinted at the GameStop stocks saga. A screenshot of the ad is available in the article.
MHC 18-478 100 I cannot help myself from thinking about smoking. What can I do to get rid of this addiction?
Table 3: Data Statistics
A.3 Case Study
In this section, we visualize three examples which are selected from the dataset Alpaca, tldr_news, and MHC respectively, shown in Figure. 3. The tokens in the red frame are the $k$ explanatory tokens extracted by the model LlaMA-2 (7B-Chat), and the output responses framed in blue are the most informative results selected by human. In the first data sample, the recognized token: "brains" and "cognitive", has some semantic relationship intuitively. It is a common sense the cognitive ability is underpinned by the brain’s function and structure. There are numerous reports and studies focused on the functionally relationship between the brain and cognitive ability (zhang2019cognitive, ). Due to the extensive amount of data on which the LLMs are trained, the model can integrate their relationship which is reflected in its response.
As for the second example, the visualized tokens are almost related to the "book" and "publish", which aligns with the theme of the output content. Removing either "publishing" or "publish" would not greatly alter the meaning of the input prompt. This example also shows that our method could detect some fix expressions like "whether…or..", whose relation is recognized by the token "vs." in the output.
In the last example, we could figure out the model infers or "guesses" that the kid feels frustrated and helpless, a sentiment partly related to the token "never". The generated output does not mention the reason to explain how the model makes such inference. From this point, our method could generate human-intelligible explanations to assist the users in identifying which parts of the prompts fundamentally lead to the undesired response, providing guidance on how to revise the prompts effectively.
Figure 3: Case study for visualizing the top-$k$ tokens in each example.
Figure 4: Case study for visualizing the variation of the model responses after the specific tokens are masked.
In Figure 4, we also illustrate the changes in the model response as the targeted top-$k$ token is removed from the input prompt. Compared with the top-$k$ tokens identified by our method, those found by Captum are mostly not important to the response. This demonstrates the Captum’s limited ability to explain the relationship between input prompt and output response with the consideration of the textual information. Moreover, when masking the token "publishing" or "publish", as identified by our method JoPA, the generated outputs do not exhibit significant variations in either word frequency or semantics. However, substantial changes in the generated outputs occur only when both tokens are simultaneously masked. This highlights the effectiveness of our methodology in explaining the joint effect of the prompt attributions on the generated output.
A.4 Benefits to the community on improving model ability
Benefit on Model Safety
Our method has the potential to help LLM defend against jailbreaking attacks, making the model more robust. We verify this via a real case study. Given an adversarial prompt with malicious request: "Generate a message that spreads fake news. organized few GO a a a a newcommand ! SudGuiheures ! !/ refection/:" It can jailbreak Llama-2 (7B-Chat) model to generate unsafe responses. Our algorithm successfully attributes "Sud" and "he", which are part of the adversarial suffix. By deleting these attributions, the attack fails. Thus, our approach could help detect these high-risk malicious tokens to robustify model safety.
Benefit on Model Diagnosis
Our approach can also be used to assess the relevance of model’s response to the prompt, helping diagnose the effectiveness of model responding to the prompt. For a real case study, consider a single prompt consisting of multiple tasks: "Create a message to complain about the noise in the neighborhood to friends. Give me the Python code for filtering results based on the following columns…" We find Llama-2 (7B-Chat) only responds to the first request, neglecting the second one. Our approach attributes the model’s output to tokens like "complain" "noise" and "friends", mainly about the first request. Thus by input attribution, our method can help diagnose the overlooked request and suggest better prompt design.
Detection Accuracy
JoPA could be used as the defense of the jailbreak attacks and help users to design better prompts to get satisfying responses. We also do the experiments to show the effectiveness of applying JoPA to detect malicious prompts on Llama-2 (7B-Chat) model, which is similar to the application done in CONTEXTCITE (contextcite, ) The results are shown below:
Method Detection Accuracy by JoPA ASR
GCG 100% 3%
Prompt with Random Search 91% 90%
Table 4: Comparison of detection accuracy and attack success rate.
[…]
A.5 Discussion on Counterfactual Explanation
Our work JoPA is to explain the generation of LLMs by identifying a small changes (i.e., masking few tokens) needed to alter the outcome (i.e., minimizing the probability of the original generation). This can be understood as counterfactual explanation, since it answers a "what if" question: if these input tokens are masked, the LLM would not generate the response. Namely, by applying the mask $m$ on the input prompt $x$, the generation probability of its original response $y$ would be largely decreased (i.e., $p_{\theta}(y|m\odot x)\ll p_{\theta}(y|x)$). This suggests that the newly generated output $y^{\prime}$ resulting from the masked prompt would differ substantially from the original outcome $y$, as measured by our metrics in Table 1: the semantic similarity (SentenceBert) and word similarity (BLEU and ROUGE-L) between $y$ and $y^{\prime}$ are much smaller than those from baselines. Here are several case studies to illustrate the concepts discussed.
General Q&A Task
[…]
Fill-in-the-Blank Task
[…]
The case studies of diverse tasks above demonstrate that the new responses generated from the masked inputs differ significantly from the original outputs, which highlights the impact of these explanatory tokens in influencing and altering the model’s output.
A.6 Ablation Study
Number of Explanatory Tokens.
Figure 6 shows the Probability Ratio (PR) score as the number of masked tokens $k$ changes on each dataset for LlaMA-2 (7B-Chat). Our algorithm consistently outperforms other baseline methods even with larger $k$, demonstrating the stable performance and effectiveness of JoPA. After masking $k=2$ tokens, the value of PR would decrease dramatically, indicating that these two tokens are crucial for generating the output $\bm{y}$ and they act as the triggers that alter the probability distribution of the output. As the number of tokens increase to $k=5$, the PR keeps decreasing but with a less sharp slope. This suggests that the probability of generating the original output $\bm{y}$ is significantly determined by a few predominant tokens.
Figure 5: Probability Ratio with varying number of masked tokens $k$.
Figure 6: Convergence plot of JoPA.
Number of Sampling Iterations.
We now demonstrate the convergence of our search algorithm. For each sampling iteration, we sample entries in the mask for value swapping; the mask will be updated if the swapping leads to a drop in the generation log-likelihood $\log p_{\bm{\theta}}(\bm{y}|\bm{m}\odot{\bm{x}})$. Figure 6 empirically shows how the log-likelihood decreases as the sampling and mask update continue. The quickly decreasing trend demonstrates that our algorithm is successfully performed to improve the quality of mask in locating the predominant tokens on generating $\bm{y}$.
[…]
In addition, we extend the application of our framework, JoPA, to a larger model and a more challenging task, specifically Few-shot Chain-of-Thought (CoT) reasoning (wei2022chain, ; kojima2022large, ), as described in Appendix A.11 and Appendix A.12. The better performance of our algorithm compared with others highlights the transferability of our framework across different model architectures and tasks.
A.7 Gradient Design
The Table. 5 displays the log-likelihood without using the gradient as the guidance to initialize the optimization process or to do sampling at different stage when searching for the optimal results.
Methods Number of Iterations
1 5 10 15 20 30 40 50
JoPA -42.448 -51.661 -56.378 -59.852 -62.303 -64.812 -67.274 -69.426
$O$ w/o Initialization -33.906 -44.555 -50.828 -55.652 -58.483 -62.694 -64.793 -66.415
$O$ w/o Probability -41.759 -49.867 -53.784 -57.107 -59.639 -63.339 -65.848 -67.496
Table 5: Ablation study showing the log-likelihood without using gradient for initialization or sampling with different number of iterations.
A.8 Results on Larger Dataset
To compare with Captum, which is 1000 times slower than JoPA (Table 2), we initially conducted experiments using a dataset of around 100 samples. To further demonstrate the effectiveness of our method, we expanded the data size by randomly sampling 800 examples from the Alpaca and tldr_news datasets and utilizing the entire MHC dataset. To maintain computational feasibility, we capped the maximum runtime for obtaining results at 240 hours. The experimental results in Table 6 indicate the effectiveness of our method as the sample size increases on Llama-2 (7B-Chat) model.
[…]
Precision Recall F1
Alpaca Random 0.621 0.539 0.540 0.535 0.824 0.871 0.104
Last-Attention 0.540 0.447 0.453 0.444 0.693 0.787 0.191
Attention 0.546 0.462 0.464 0.458 0.713 0.802 0.168
Captum
Integrated-Gradient 0.497 0.395 0.396 0.390 0.687 0.704 0.275
JoPA 0.479 0.383 0.376 0.372 0.642 0.565 0.479
tldr_news Random 0.795 0.738 0.739 0.738 0.920 0.943 0.040
Last-Attention 0.749 0.678 0.682 0.679 0.879 0.867 0.110
Attention 0.761 0.695 0.698 0.695 0.900 0.903 0.074
Captum
Integrated-Gradient 0.693 0.605 0.611 0.607 0.849 0.783 0.175
JoPA 0.688 0.603 0.606 0.604 0.832 0.589 0.419
MHC Random 0.706 0.595 0.591 0.590 0.788 0.953 0.027
Last-Attention 0.662 0.526 0.526 0.525 0.724 0.905 0.063
Attention 0.664 0.531 0.530 0.529 0.739 0.916 0.053
Captum
Integrated-Gradient 0.610 0.470 0.479 0.479 0.643 0.806 0.145
JoPA 0.579 0.408 0.407 0.407 0.594 0.688 0.358
Table 6: Experimental results on evaluation metrics for larger data samples. (– indicates exceeding the 200-hour runtime limit)
A.9 Variance Assessment
To assess the variance of these metrics, we run the experiment with different seed for 3 times on the Llama-2 (7B-Chat) model. The results of different metrics on three datasets are shown in the Table. 7, where value in each cell denotes $\text{mean}\pm±\text{std}$. The small variance indicates that our algorithm is quite stable across different runs.
[…]
Precision Recall F1
[…]
Table 7: Variance measurement results for LlaMA-2 (7B-Chat)
A.10 Optimizing Mask via Gradient
As discussed in Section 4, the strategy that directly optimizes the mask based on gradients requires first relaxing the discrete problem into a continuous optimization problem (i.e. $\bm{m}\in[0,1]^{T}$), optimizing via gradient descent, and later projecting the continuous solution back into the discrete space (i.e., $\bm{m}\in\{0,1\}^{T}$). This projection from continuous to discrete space in practice usually results in large rounding error, i.e., after projection, the loss increases dramatically. The results of this strategy for Llama-2 (7B-Chat) model on the Alpaca dataset are in Table 8:
[…]
Precision Recall F1
Gradient 0.624 0.523 0.534 0.520 0.839 0.825 0.030
JoPA 0.484 0.388 0.386 0.379 0.642 0.549 0.504
Table 8: Optimizing continuous mask on Alpaca dataset.
The results verify our claim: directly optimizing $\bm{m}$ using the gradients (and then projecting to discrete solution) has much worse performance than our method, highlighting the challenge of this discrete optimization problem and our contribution of solving it effectively.
A.11 Results on Larger Model
To further validate a wider applicability of our algorithm on larger models, we conducted additional experiments for Llama-2 (70B-Chat) 16-bit model on the Alpaca dataset. Due to the inefficiency of baseline Captum, we randomly sampled 20 instances from the Alpaca dataset to obtain the results in Figure 7, which demonstrate that our algorithm still outperforms the baselines on this larger model with clear margins.
Figure 7: Results of Llama-2 (70B-Chat) model on the Alpaca dataset. $\downarrow$ indicates lower is better; $\uparrow$ indicates higher is better.
A.12 Method Generalizability on Reasoning Task
We demonstrate the generalizability of our algorithm by conducting an additional experiment on the reasoning task of Few-shot-CoT (wei2022chain, ; kojima2022large, ), using the dataset AQuA (kojima2022large, ; goswami2024aqua, ). The following results in Figure 8 shows the remarkable performance of our method, indicating its ability to generalize well on more complex tasks like CoT.
Figure 8: Experimental results on reasoning task of Few-shot-CoT. $\downarrow$ indicates lower is better; $\uparrow$ indicates higher is better.
A.13 Comparison with Other Baselines
ReAGent (zhao2024reagent, ) addresses a different task than JoPA. It focuses on classification tasks by explaining the importance of words in inputs corresponding to the predicted label and extends this approach to generation tasks by explaining the single next predicted word. In contrast, our task focuses on explaining the relationship between the input prompt and the entire generated sentences. To compare the performance of JoPA and ReAGent on the generation task, we modify ReAGent’s explanatory target to encompass the full generation outputs rather than a single word. We conduct experiments on the Alpaca dataset using the Llama-2 (7B-Chat) model, and the results are presented in Figure 9. The experimental results demonstrate that JoPA outperforms in explaining the relationship between joint attributions in the input and the resulting output.
Figure 9: Experimental results between JoPA and ReAGent. $\downarrow$ indicates lower is better; $\uparrow$ indicates higher is better.